# On the complementation of spaces of $\mathcal{I}$-null sequences

### v1

@@ Abstract.
We study the complementation (in $\ell_{\infty}$) of the Banach space $c_{0,\mathcal{I}}$, consisting of all bounded sequences $(x_{n})$ that $\mathcal{I}$-converge to 00, endowed with the supremum norm, where $\mathcal{I}$ is an ideal of subsets of $\mathbb{N}$. We show that the complementation of these spaces is related to a condition requiring that the ideal is the intersection of a countable family of maximal ideals, which we refer to as $\omega$-maximal ideals. We prove that if $c_{0,\mathcal{I}}$ admits a projection satisfying a certain condition, then $\mathcal{I}$ must be a special type of $\omega$-maximal ideal. Additionally, we characterize when the quotient space $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$ is finite-dimensional for two ideals $\mathcal{I}\subsetneq\mathcal{J}$.
@@ 1. Introduction
We call a proper ideal $\mathcal{I}$ complemented if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. Leonetti [18] proved that any meager ideal is not complemented, where a meager ideal refers to an ideal that is meager as a subset of the Cantor cube $\{0,1\}^{\mathbb{N}}$, identified via characteristic functions. The key element of his argument is the existence of uncountable families of subsets in $\mathcal{P}(\mathbb{N})\setminus\mathcal{I}$ such $A\cap B\in\mathcal{I}$ for any $A,B$ in the family (the so-called $\mathcal{I}$-AD families). We show that if $\mathcal{I}$ is complemented, then any $\mathcal{I}$-AD family is at most countable and therefore $\mathcal{I}$ is not meager and does not have the Baire property as a subset of $\{0,1\}^{\mathbb{N}}$.
On the other hand, Kania [14] observed that the intersection of a finite collection of maximal ideals is complemented (for a proof, see [23]). We call an ideal $\omega$-maximal if it can be written as a countable intersection of maximal ideals. We extend Kania’s result to certain special $\omega$-maximal ideals.
@@
In the last section, we discuss the problem of the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$. One of our main results is that any strongly $\omega$-maximal ideal is complemented (see Theorem 5.8). We provide two different proofs of this result. The first one (see Theorem 5.2) is shorter but considerably less informative, as it relies on a classical theorem of Lindenstrauss stating that a closed subspace of $\ell_{\infty}$ is complemented if and only if it is isomorphic to $\ell_{\infty}$. Furthermore, we characterize strongly $\omega$-maximal ideals in terms of the type of projections on their associated space $c_{0,\mathcal{I}}$ (see Theorem 5.20).
In addition, we show that if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$ by a projection satisfying an extra condition, then $\mathcal{I}$ must be an $\omega$-maximal ideal (see Theorem 5.18). Whether the converse holds—that is, whether every $\omega$-maximal ideal is complemented—remains an open question.
@@ 2. Preliminaries
We will use standard terminology and notation for Banach lattices and Banach space theory. For unexplained definitions and notations, we refer to [1, 24]. The scalar field is denoted by $\mathbb{K}$. All Banach lattices analyzed here are assumed to be real. However, our results can be extended to complex Banach lattices in the usual manner [24, Chapter 2, p. 133]. If $X$ and $Y$ are isomorphic Banach spaces, we write $X\sim Y$. If $E$ is a closed subspace of a Banach space $X$, we say that $E$ is complemented in $X$ if there is a continuous onto operator $P\colon X\to E$ such that $P^{2}=P$, or equivalently, there is a closed subspace $W$ of $X$ such that $X=E\oplus W$. In addition, if $E$ and $X$ are Banach lattices, $P$ is called positive if $Px\geq{\bf 0}$ for all $x\geq{\bf 0}$.
@@ 3. $\omega$-maximal ideals
Since we are always assuming that ${\sf Fin}$ is contained in any ideal under consideration, every maximal ideal extending a given ideal is necessarily non-principal. It is worth keeping in mind that the intersection of less than $2^{\aleph_{0}}$ maximal ideals on $\mathbb{N}$ does not have the Baire property, and hence it is non-meager ([22], [26, Proposition 23]). Our main interest will be on $\omega$-maximal ideals since they are related to the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$.
@@
[…]
@@
For each collection of maximal ideals $\mathfrak{E}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ we define a family of infinite subsets of $\mathbb{N}$ which will be very helpfull for what follows:
@@
For the last claim, suppose $\mathfrak{E}_{1}$ and $\mathfrak{E}_{2}$ are two discrete countable collections of maximal ideals such that $\bigcap\mathfrak{E}_{1}=\bigcap\mathfrak{E}_{2}$. Let $\{J_{n}:n\in\mathbb{N}\}$ be an enumeration (without repetitions) of $\mathfrak{E}_{1}\cup\mathfrak{E}_{2}$. Since $\bigcap\mathfrak{E}_{1}=\bigcap(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$, $\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ has a unique maximal element $M$. Suppose there is $n_{0}$ such that $J_{n_{0}}\in\mathfrak{E}_{2}\setminus\mathfrak{E}_{1}$. Let $L=\{n\in\mathbb{N}:J_{n}\in\mathfrak{E}_{2}\}$ and notice that $L\in\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ is maximal by Lemma 3.3 and thus $n_{0}\in M$. On the other hand, $\bigcap\{J_{n}:n\in M\setminus\{n_{0}\}\}=\bigcap\mathfrak{E}_{2}\subseteq J_{n_{0}}$ and, by Lemma 3.2, $M$ is not discrete, a contradiction. ∎
@@
Suppose $\mathfrak{E}^{*}$ is scattered. Since the collection of isolated points of $\mathfrak{E}^{*}$ is discrete and dense in $\mathfrak{E}^{*}$ (see [25, p. 150]), there is $M\in\mathcal{DI}(\mathfrak{E})$ such that $\{\mathcal{I}^{*}:\,n\in M\}$ is dense in $\mathfrak{E}^{*}$, then easily $M$ is maximal in $\mathcal{DI}(\mathfrak{E})$. Thus by Theorem 3.4, $\mathcal{I}$ is strongly $\omega$-maximal. The converse is obvious. ∎
@@
[…]
@@
Let $\mathcal{F}$ be a filter on $\mathbb{N}$ (containing all co-finite sets). Define a topology $\tau_{\mathcal{F}}$ over ${\mathbb{N}}^{<\omega}$ by letting a subset $U$ of ${\mathbb{N}}^{<\omega}$ be open, if $\{n\in\mathbb{N}:\;s\widehat{\;}n\in U\}\in\mathcal{F}$ for all $s\in U$. Then $({\mathbb{N}}^{<\omega},\tau_{\mathcal{F}})$ is Hausdorff, zero-dimensional and without isolated points. Moreover, when $\mathcal{F}$ is a non-principal ultrafilter, $({\mathbb{N}}^{<\omega},\tau_{\mathcal{F}})$ is extremally disconnected, i.e., the closure of an open set is open (see [20] and [6]). Since this space is zero-dimensional and has no isolated points, for every discrete $D\subseteq{\mathbb{N}}^{<\omega}$ there is $s\in{\mathbb{N}}^{<\omega}\setminus D$ such that $D\cup\{s\}$ is still discrete, i.e., there are no maximal discrete subsets of ${\mathbb{N}}^{<\omega}$. It is a classical fact that every countable extremally disconnected Hausdorff space can be embedded in $\beta\mathbb{N}$ (see, for instance, [27, Theorem 1.4.7]). Therefore, if $\mathfrak{E}\subseteq\beta\mathbb{N}$ is homeomorphic to our space, then $\mathfrak{E}$ has no maximal discrete subsets, that is =$\mathcal{I}=\bigcap\{\mathcal{U}^{*}:\mathcal{U}\in\mathfrak{E}\}$ is an $\omega$-maximal ideal which is not strongly $\omega$-maximal (by Theorem 3.4).
@@ 4. Dimension of $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$
Conversely, suppose $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ and let $\{\mathcal{L}_{1},\ldots,\mathcal{L}_{k}\}$ be a family of maximal ideals on $\mathbb{N}$ with $\mathcal{J}\not\subset\mathcal{L}_{i}$ for all $1\leq i\leq k$ and $\mathcal{I}=(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$. By Theorem 3.1, there is a partition $\{D_{1},\ldots,D_{k}\}$ of $\mathbb{N}$ such that $D_{i}\in\mathcal{L}^{*}_{i}\cap\mathcal{L}_{j}$ for all $j,i$ in $\{1,\ldots,k\}$ with $j\neq i$. For each $1\leq i\leq k$, let $X_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$. Set $A_{i}=X_{i}\cap D_{i}$ for $i\in\{1,\ldots,k\}$. Then $A_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$ for all $i$. We will prove that $\{A_{1},\ldots,A_{k}\}$ is a $(\mathcal{I},\mathcal{J},k)$-family. Clearly, the conditions (1) and (2) of Definition 4.2 hold. It remains to check conditions (3) and (4):
@@
[…]
@@
[…]
@@
    Since $A_{i}\cap A_{j}\in\mathcal{I}$ for all $i,j\in\{1,\ldots,k\}$ and $i\neq j$, by a standard procedure we can make them pairwise disjoint and thus we may assume that $A_{i}\cap A_{j}=\emptyset$ for each $i,j\in\{1,\ldots,k\}$ with $i\neq j$.
@@
5. When is $c_{0,\mathcal{I}}$ complemented in $\ell_{\infty}$?
In this section we address the problem of when an ideal $\mathcal{I}$ is complemented, that is, when $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. We recall that a classical theorem of Lindenstrauss [19], which states that a closed subspace $X\subseteq\ell_{\infty}$ is complemented if and only if $X$ is isomorphic to $\ell_{\infty}$. As it was noticed by Kania in [14], if $\mathcal{I}$ is the intersection of finitely many maximal ideals, then $\mathcal{I}$ is complemented (for a proof see [23, Proposisition 5.20]). Therefore, if $\mathcal{I}$ is the intersection of finitely many maximal ideals on an infinite set $A$, then $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}(A)$ (and hence, isomorphic to $\ell_{\infty}$).
[…]
We begin by stating a lemma which is a well-known consequence of Theorem 2.1 (for a proof, see, for instance, [18]).
Lemma 5.1.
On the other hand, $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})\leq\aleph_{0}$ for every $\omega$-maximal ideal $\mathcal{I}$, this was implicitly shown by Plewik [22] and, for the sake of completion, we include a direct proof. Let $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$ be an $\omega$-maximal ideal and suppose $\mathcal{B}\subseteq\mathcal{I}^{+}$ is an uncountable $\mathcal{I}$-AD family. Then, there is $n$ such that $|\mathcal{B}\cap\mathcal{I}_{n}^{+}|\geq 2$. If $A,B\in\mathcal{B}\cap\mathcal{I}_{n}^{+}$ are two different sets, then $A\cap B\in\mathcal{I}\subseteq\mathcal{I}_{n}$, which contradicts that $\mathcal{I}_{n}$ is maximal. Additionally, any $\omega$-maximal ideal is not measurable with respect to the usual product measure on $2^{\mathbb{N}}$ (see [15, section 4.1]).
[…]

### v2

@@ Abstract.
We study the complementation (in $\ell_{\infty}$) of the Banach space $c_{0,\mathcal{I}}$, consisting of all bounded sequences $(x_{n})$ that $\mathcal{I}$-converge to 0, endowed with the supremum norm, where $\mathcal{I}$ is an ideal of subsets of $\mathbb{N}$. We show that the complementation of these spaces is related to a condition requiring that the ideal is the intersection of a countable family of maximal ideals, which we refer to as $\omega$-maximal ideals. We prove that $\mathcal{I}$ is $\omega$-maximal exactly when $c_{0,\mathcal{I}}$ is the kernel of an operator from $\ell_{\infty}$ to itself satisfying certain property. In addition, we show that the existence of a Banach lattice embedding from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$ is equivalent to $\mathcal{I}$ being a $\omega$-maximal ideal. Finally, we show that $\mathcal{I}$ is a strongly $\omega$-maximal ideal if and only if the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$. Additionally, we characterize when the quotient space $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$ is finite-dimensional for two ideals $\mathcal{I}\subsetneq\mathcal{J}$.
@@ 1. Introduction
We call a proper ideal $\mathcal{I}$ complemented if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. There has been some recent interest in this notion [13, 15, 20]. Leonetti [20] proved that any meager ideal is not complemented, where a meager ideal refers to an ideal that is meager as a subset of the Cantor cube $\{0,1\}^{\mathbb{N}}$, identified via characteristic functions. The key element of his argument is the existence of uncountable families of subsets in $\mathcal{P}(\mathbb{N})\setminus\mathcal{I}$ such $A\cap B\in\mathcal{I}$ for any $A,B$ in the family (the so-called $\mathcal{I}$-AD families). We show that if $\mathcal{I}$ is complemented, then any $\mathcal{I}$-AD family is at most countable and therefore $\mathcal{I}$ is not meager and does not have the Baire property as a subset of $\{0,1\}^{\mathbb{N}}$.
On the other hand, Kania [15] observed that the intersection of a finite collection of maximal ideals is complemented (for a proof, see [25]). We call an ideal $\omega$-maximal if it can be written as a countable intersection of maximal ideals. We extend Kania’s result to certain special $\omega$-maximal ideals. Finally, continuing this approach, Hrušák-Sáenz [13] found very recently a characterization of complemented ideals in terms of the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$. Namely, $\mathcal{I}$ is complemented if and only if the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}$.
Motivated by the discussion above, our work focuses on analyzing the structural properties of complemented ideals.
@@
In Section 5, we discuss properties of complemented ideals. We show that the sum and the countable intersection of complemented ideals is complemented. So, in particular, every $\omega$-maximal ideal is complemented. We also prove that any strongly $\omega$-maximal ideal is complemented by showing an explicit projection (see Theorem 5.14). Furthermore, we characterize strongly $\omega$-maximal ideals in terms of the type of projections on their associated space $c_{0,\mathcal{I}}$ (see Theorem 5.16). In addition, we show that the existence of a Banach lattice embedding from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$ to $\ell_{\infty}$ is equivalent to $\mathcal{I}$ being a $\omega$-maximal ideal (see Theorem 5.11). Finally, we show that $\mathcal{I}$ is a strongly $\omega$-maximal ideal if and only if the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$ (see Theorem 5.19).
@@ 2. Preliminaries
We will use standard terminology and notation for Banach lattices and Banach space theory. For unexplained definitions and notations, we refer to [1, 26]. The scalar field is denoted by $\mathbb{K}$. All Banach lattices analyzed here are assumed to be real. However, our results can be extended to complex Banach lattices in the usual manner [26, Chapter 2, p. 133]. If $X$ and $Y$ are isomorphic Banach spaces, we write $X\sim Y$. If $E$ is a closed subspace of a Banach space $X$, we say that $E$ is complemented in $X$ if there is a continuous onto operator $P\colon X\to E$ such that $P^{2}=P$, or equivalently, there is a closed subspace $W$ of $X$ such that $X=E\oplus W$. In addition, if $E$ and $X$ are Banach lattices, $P$ is called positive if $Px\geq{\bf 0}$ for all $x\geq{\bf 0}$. Also, a bounded linear operator $T\colon E\to X$ is called Banach lattice homomorphism if $T(x\vee y)=Tx\vee Ty$ for each $x,y\in E$.
@@ 3. $\omega$-maximal ideals
Since we are always assuming that $\sf Fin$} is contained in any ideal under consideration, every maximal ideal extending a given ideal is necessarily non-principal. It is worth keeping in mind that the intersection of less than $2^{\aleph_{0}}$ maximal ideals on $\mathbb{N}$ does not have the Baire property, and hence it is non-meager ([24], [28, Proposition 23]). Our main interest will be on $\omega$-maximal ideals since they are related to the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$.
@@
    $\mathcal{I}$ is a limit point of $\mathfrak{D}$ if and only if $\bigcap\mathfrak{D}=\bigcap(\mathfrak{D}\setminus\{\mathcal{I}\})$. Consequently, $\mathfrak{D}$ is discrete if and only if $\bigcap(\mathfrak{D}\setminus\{\mathcal{K}\})\not\subseteq\mathcal{K}$ for each $\mathcal{K}\in\mathfrak{D}$.
@@
For each collection of maximal ideals $\mathfrak{E}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ we define a family of infinite subsets of $\mathbb{N}$ which will be very helpful for what follows:
@@
For the last claim, suppose $\mathfrak{E}_{1}$ and $\mathfrak{E}_{2}$ are two discrete countable collections of maximal ideals such that $\bigcap\mathfrak{E}_{1}=\bigcap\mathfrak{E}_{2}$. Let $\{J_{n}:n\in\mathbb{N}\}$ be an enumeration (without repetitions) of $\mathfrak{E}_{1}\cup\mathfrak{E}_{2}$. Since $\bigcap\mathfrak{E}_{1}=\bigcap(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$, $\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ has a unique maximal element $M$. Suppose there is $n_{0}$ such that $J_{n_{0}}\in\mathfrak{E}_{2}\setminus\mathfrak{E}_{1}$. Let $L=\{n\in\mathbb{N}:J_{n}\in\mathfrak{E}_{2}\}$ and notice that $L\in\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ is maximal by Lemma 3.3 and thus $n_{0}\in M$. On the other hand, $\bigcap\{J_{n}:n\in M\setminus\{n_{0}\}\}=\bigcap\mathfrak{E}_{2}\subseteq J_{n_{0}}$ and, by Proposition 3.2, $M$ is not discrete, a contradiction. ∎
@@
Suppose $\mathfrak{E}^{*}$ is scattered. Since the collection of isolated points of $\mathfrak{E}^{*}$ is discrete and dense in $\mathfrak{E}^{*}$ (see [27, p. 150]), there is $M\in\mathcal{DI}(\mathfrak{E})$ such that $\{\mathcal{I}^{*}:\,n\in M\}$ is dense in $\mathfrak{E}^{*}$, then easily $M$ is maximal in $\mathcal{DI}(\mathfrak{E})$. Thus by Theorem 3.4, $\mathcal{I}$ is strongly $\omega$-maximal. The converse obviously follows from Theorem 3.4. ∎
@@
Following [16], a collection of non principal ultrafilters $\mathfrak{M}$ is said minimal, if $\bigcap\mathfrak{M}\neq\bigcap(\mathfrak{M}\setminus\{\mathcal{F}\})$ for every $\mathcal{F}\in\mathfrak{M}$. A filter $\mathcal{F}$ has a minimal representation if $\mathcal{F}=\bigcap\mathfrak{M}$ for some minimal collection $\mathfrak{M}$ of ultrafilters. Lemma 3.2 says that $\mathfrak{M}$ is minimal if and only if it is a discrete subset of $\beta\mathbb{N}$. And, from Theorem 3.5 it follows that a filter $\mathcal{F}$ has a countable minimal representation if and only if $\mathcal{F}^{*}$ is strongly $\omega$-maximal ideal. They also showed that whenever a filter admits a countable minimal representation, it is unique. This also follows from Theorem 3.4.
@@
Let $\mathcal{F}$ be a filter on $\mathbb{N}$ (containing all co-finite sets). Define a topology $\tau_{\mathcal{F}}$ over ${\mathbb{N}}^{<\omega}$ by letting a subset $U$ of ${\mathbb{N}}^{<\omega}$ be open, if $\{n\in\mathbb{N}:\;s\widehat{\;}n\in U\}\in\mathcal{F}$ for all $s\in U$. Then $({\mathbb{N}}^{<\omega},\tau_{\mathcal{F}})$ is Hausdorff, zero-dimensional and without isolated points. Moreover, when $\mathcal{F}$ is a non-principal ultrafilter, $({\mathbb{N}}^{<\omega},\tau_{\mathcal{F}})$ is extremally disconnected, i.e., the closure of an open set is open (see [22] and [6]). Since this space is zero-dimensional and has no isolated points, for every discrete $D\subseteq{\mathbb{N}}^{<\omega}$ there is $s\in{\mathbb{N}}^{<\omega}\setminus D$ such that $D\cup\{s\}$ is still discrete, i.e., there are no maximal discrete subsets of ${\mathbb{N}}^{<\omega}$. It is a classical fact that every countable extremally disconnected Hausdorff space can be embedded in $\beta\mathbb{N}$ (see, for instance, [29, Theorem 1.4.7]). Therefore, if $\mathfrak{E}\subseteq\beta\mathbb{N}$ is homeomorphic to our space, then $\mathfrak{E}$ has no maximal discrete subsets, that is $\mathcal{I}=\bigcap\{\mathcal{U}^{*}:\mathcal{U}\in\mathfrak{E}\}$ is an $\omega$-maximal ideal which is not strongly $\omega$-maximal (by Theorem 3.4).
@@ 4. Dimension of $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$
Conversely, suppose $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ and let $\{\mathcal{L}_{1},\ldots,\mathcal{L}_{k}\}$ be a family of maximal ideals on $\mathbb{N}$ with $\mathcal{J}\not\subset\mathcal{L}_{i}$ for all $1\leq i\leq k$ and $\mathcal{I}=(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$. By Proposition 3.1, there is a partition $\{D_{1},\ldots,D_{k}\}$ of $\mathbb{N}$ such that $D_{i}\in\mathcal{L}^{*}_{i}\cap\mathcal{L}_{j}$ for all $j,i$ in $\{1,\ldots,k\}$ with $j\neq i$. For each $1\leq i\leq k$, let $X_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$. Set $A_{i}=X_{i}\cap D_{i}$ for $i\in\{1,\ldots,k\}$. Then $A_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$ for all $i$. We will prove that $\{A_{1},\ldots,A_{k}\}$ is a $(\mathcal{I},\mathcal{J},k)$-family. Clearly, the conditions (1) and (2) of Definition 4.2 hold. It remains to check conditions (3) and (4):
@@
Let $A_{1},\ldots,A_{m}\in\mathcal{A}$ be such that $\sum_{j=1}^{m}a_{j}(\chi_{A_{j}}+c_{0,\mathcal{I}})={\bf 0}$ for some $\{a_{j}\,\colon\,1\leq j\leq m\}\subset\mathbb{K}$. Let $D_{1}=A_{1}$ and $D_{j}=A_{j}\setminus(A_{1}\cup\cdots\cup A_{j-1})$ for $2\leq j\leq m$. Since $A_{i}\cap A_{j}\in\mathcal{I}$ for all $i,j\in\{1,\ldots,k\}$ with $i\neq j$, we have $D_{j}\in\mathcal{J}\cap\mathcal{I}^{+}$ and $\chi_{A_{j}}+c_{0,\mathcal{I}}=\chi_{D_{j}}+c_{0,\mathcal{I}}$ for all $j\in\{1,\ldots,m\}$. So, $\sum_{j=1}^{m}a_{j}(\chi_{D_{j}}+c_{0,\mathcal{I}})={\bf 0}$. Thus, $a_{j}=0$ for all $1\leq j\leq m$. ∎
@@
Suppose that $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=k$. By [26, Corollary 1, p. 70], there is an order isomorphism $\phi\colon c_{0,\mathcal{J}}/c_{0,\mathcal{I}}\to\mathbb{R}^{k}$. Let $\Lambda\colon c_{0,\mathcal{J}}\to\mathbb{R}^{k}$ be given by $\Lambda({\bf y})=\phi({\bf y}+c_{0,\mathcal{I}})$ if ${\bf y}\in c_{0,\mathcal{J}}$. Notice that $\Lambda$ is an onto Banach lattice homomorphism. For each $j\in\{1,\ldots,k\}$, let ${\bf y}_{j}\in c_{0,\mathcal{J}}\setminus c_{0,\mathcal{I}}$ be such that $\Lambda({\bf y}_{j})=e_{j}$. Fix $j\in\{1,\ldots,k\}$ and let $\varepsilon_{j}>0$ be such that $A_{j}:=A(\varepsilon_{j},{\bf y}_{j})\not\in\mathcal{I}$. So, $\varepsilon_{j}\chi_{A_{j}}\leq{\bf y}_{j}$. Since $\Lambda$ is order preserving, we conclude that $\Lambda(\chi_{A_{j}})=a_{j}e_{j}$ for some $a_{j}>0$. Notice that if $i,j\in\{1,\ldots,k\}$ and $i\neq j$, we have $A_{i}\cap A_{j}\in\mathcal{I}$ because of $\Lambda(\chi_{A_{i}\cap A_{j}})=\Lambda(\chi_{A_{i}}\wedge\chi_{A_{j}})=a_{i}e_{i}\wedge a_{j}e_{j}={\bf 0}$. Thus, $\{A_{1},\ldots,A_{k}\}$ is a $\mathcal{I}$-AD family. As in the proof of Lemma 4.5, we will assume that the family is pairwise disjoint. From Lemma 4.5 we obtain that $\{\chi_{A_{j}}+c_{0,\mathcal{I}}\,\colon\,1\leq j\leq k\}$ is a basis for $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$. We claim that $\{A_{1},\ldots,A_{k}\}$ is a $(\mathcal{I},\mathcal{J},k)$-family:
@@
    $\{A_{j}:1\leq j\leq k\}$ is pairwise disjoint by construction.
@@
5. Structural properties of complemented ideals
In this section, we study the structural properties of complemented ideals $\mathcal{I}$, that is, ideals for which the Banach space $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. We recall that a classical theorem of Lindenstrauss [21] says that a closed subspace $X\subseteq\ell_{\infty}$ is complemented if and only if $X$ is isomorphic to $\ell_{\infty}$. As noted by Kania in [15], the intersection of finitely many maximal ideals is complemented (for a proof, see [25, Proposisition 5.20]). Therefore, if $\mathcal{I}$ is such an ideal, then $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$. Recently, Hrušák and Sáenz characterized the complementation of $c_{0,\mathcal{I}}$ in terms of the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ [13]. Before stating their result, we need to introduce some notation.
For an ideal $\mathcal{I}$, let $K_{\mathcal{I}}=\{p\in\beta\mathbb{N}\,\colon\,\mathcal{I}^{*}\subseteq p\}$. It is known that $K_{\mathcal{I}}$ is a compact Hausdorff space, since it is a closed subset of $\beta\mathbb{N}$. Moreover, the map
$\Psi\colon\ell_{\infty}/c_{0,\mathcal{I}}$ $\to C(K_{\mathcal{I}})$
${\bf x}+c_{0,\mathcal{I}}$ $\mapsto\Psi({\bf x}+c_{0\mathcal{I}})\,\colon\,p\in K_{\mathcal{I}}\mapsto p-\lim{\bf x}$
is a Banach lattice isometry. Therefore, we have the following:
Theorem 5.1.
(Kania [15]) $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isometric to $C(K_{\mathcal{I}})$.
Theorem 5.2.
(Hrušák and Sáenz [13]). Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. Then, the following statements are equivalent:
  1. (1)
    $\mathcal{I}$ is complemented.
  2. (2)
    $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}$.
  3. (3)
    $\ell_{\infty}/c_{0,\mathcal{I}}$ is isometric to a subspace of $\ell_{\infty}$.
  4. (4)
    $C(K_{\mathcal{I}})$ is isomorphic to a subspace of $\ell_{\infty}$.
Corollary 5.3.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. Then, $\mathcal{I}$ is complemented if and only if $c_{0,\mathcal{I}}$ is the kernel of a continuous operator from $\ell_{\infty}$ to $\ell_{\infty}$.
Proof.
If $\mathcal{I}$ is complemented, then $c_{0,\mathcal{I}}$ is a kernel of an operator from $\ell_{\infty}$ to $\ell_{\infty}$. Conversely, let $Q\colon\ell_{\infty}\to\ell_{\infty}$ be a continuous operator such that $\ker Q=c_{0,\mathcal{I}}$. Then, the map $\hat{Q}\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}$ defined by ${\bf x}+c_{0,\mathcal{I}}\mapsto Q({\bf x})$, is an embedding. By Theorem 5.2, $\mathcal{I}$ is complemented. ∎
Recall that if $(X_{j})_{j\in\mathbb{N}}$ is a sequence of Banach spaces, the space $\ell_{\infty}((X_{j})_{j\in\mathbb{N}})$ denotes their $\ell_{\infty}$-sum, i.e., the Banach space of all bounded sequences $(x_{j})\in\prod_{j}X_{j}$, equipped with the norm $\|\cdot\|$ given by $\|(x_{j})\|=\sup_{j}\|x_{j}\|$.
Remark 5.4.
Let $(X_{n})_{n\in\mathbb{N}}$ and $(Y_{n})_{n\in\mathbb{N}}$ be sequences of Banach spaces (lattices) and assume that $T_{n}\colon X_{n}\to Y_{n}$ is an isomorphism (Banach lattice isomorphism) for each $n\in\mathbb{N}$. Then, $\ell_{\infty}((X_{n})_{n\in\mathbb{N}})$ is isomorphic (Banach lattice isomorphic) to $\ell_{\infty}((Y_{n})_{n\in\mathbb{N}})$ via the map $T\colon(x_{n})\mapsto(T_{n}(x_{n}))$. In particular, if $\{K_{n}\,\colon\,n\in\mathbb{N}\}$ is a partition of $\mathbb{N}$, then $\ell_{\infty}((\ell_{\infty}(K_{m}))_{m\in\mathbb{N}})$ is isomorphic to $\ell_{\infty}$, since $\ell_{\infty}(K_{n})\sim\ell_{\infty}$ for each $n\in\mathbb{N}$ and $\ell_{\infty}(\ell_{\infty})\sim\ell_{\infty}$.